# Conditional Information Bottleneck for Multimodal Fusion: Overcoming Shortcut Learning in Sarcasm Detection

### v1

@@ Abstract
1School of Artificial Intelligence, Beijing Normal University, Beijing, China
2IEIT SYSTEMS Co., Ltd., Beijing, China
wangyihua@mail.bnu.edu.cn, jiaqi01@ieisystem.com,
xucong@ieisystem.com, chenfeiyuchina@mail.bnu.edu.cn,
zxyuhan@mail.bnu.edu.cn, zhanghaotian@mail.bnu.edu.cn
1 Introduction
Multimodal sentiment analysis (MSA) [50] integrates various modalities to interpret human emotions, sentiments, and opinions. In MSA tasks, multimodal sarcasm detection [3] is particularly challenging due to the subtle contrast between surface meaning and underlying intent, which may convey humor, ridicule, or contempt [24]. See Figure 1, Sheldon’s words express surprise, but his facial expression shows disgust, and his tone is mocking, conveying sarcasm. Despite advances in large language models, their performance remains capped by supervised baselines, falling short of human-level understanding [22].
@@ 1 Introduction
Numerous studies have explored multimodal sarcasm detection (MSD) from various perspectives. Many focus on the inherent characteristics of sarcasm or are specifically designed for prior information: such as contrasting emotional polarity [30, 44], or incorporate external knowledge [49], character profiles [26], emotional cues [14], and fine-grained visual features [40].
As shown by issues 1, 2, and 3 in Figure 1, models tend to learn specific shortcuts. Those shortcuts [15] often fail to generalize to more diverse conditions, typically in real-world scenarios. We identify three major shortcut issues in multimodal sarcasm detection research with the representative dataset MUStARD++ [35]. (1) Heuristics in Character Role Labeling: In sarcasm datasets, some characters exhibit a speaking style that favors sarcasm, such as Sheldon in the TV show THE BIG BANG THEORY. Introducing these character labels during training causes the model to form a bias toward sarcasm associated with a specific person. (2) Label Leakage from Canned Laughter: Canned laughter (artificial background laughter) often causes label leakage. A review of MUStARD++ reveals that canned laughter frequently follows sarcastic utterances, but rarely non-sarcastic ones. Thus, the model tends to classify samples with canned laughter as sarcasm. (3) Shortcuts of Emotional Inconsistency: The dataset provides implicit/explicit emotion labels, and their contrast serves as a strong indicator of sarcasm. Some models are designed to learn implicit/explicit emotions rather than true sarcasm, limiting their ability to detect sarcasm due to over-reliance on inconsistency.
The key to understanding the intended complex information and genuinely improving performance is addressing the core multimodal fusion problem, rather than designing modules based on shortcut learning. Various fusion methods [35, 53, 54] do not result in significant information gains, some even reducing task performance, highlighting their poor generalization and adaptability to sarcasm tasks. Higher modal complementarity leads to poorer robustness when certain modes are absent or misaligned, and noise has a greater impact on modes with high complementarity [25]. Multimodal sarcasm detection involves heterogeneous information with complementary but also irrelevant or misleading redundancy. Therefore, this underscores the need for an efficient fusion strategy that integrates complementary information while mitigating redundancy.
@@ 2 Related Work
Multimodal Sarcasm Detection has evolved from text [16] and image-text [45] to video clips, which better reflect real-life scenarios. Many methods leverage sarcasm traits to aid recognition, such as incongruity [30]. Song et al. [38] propose an utterance-level attention and incongruity learning network to capture incongruity representations in sarcastic expressions. Since sarcasm is closely linked to emotions and sentiments, an adaptive representation learning model [53] is proposed based on an expert network to analyze the emotion and sarcasm tasks jointly. Chauhan et al. [5] design a multitasking collaborative framework for cross-training and sharing attention between sarcasm, implicit/explicit emotions, and implicit/explicit sentiments. Additionally, Zhang et al. [54] employ a dual-gating network and three separate layers to achieve cross-modal interactive learning. Other research attempts to leverage various external information. Studies such as [27] and [49] have utilized tools like KnowleNet to implement knowledge fusion networks for MSD. Moreover, works like [43], [40], and [34] enhance sarcasm recognition by incorporating conversational behavioral cues, fine-grained eye gaze information, or deep visual features acquired through self-conditioning ConvNet, respectively. While [14, 21] rely on pre-training with explicit and implicit emotions, the best results from [26] similarly through character labeling. Besides, some novel approaches [29, 41] suggest that quantum mechanics are well-suited for capturing the complexity and uncertainty in sarcasm through a quantum probability-driven model or quantum fuzzy neural networks.
@@
Information Bottleneck (IB) based approaches [37, 1] have demonstrated their effectiveness in various tasks such as cross-modal clustering [48] and representation learning [51, 9], and have also been widely applied in multimodal fusion. For example, Zhang et al. [52] improve multimodal sentiment analysis performance by applying IB constraints to pairs of modalities. Mai et al. [31] ensure that latent modal representations can effectively handle the target task by introducing IB constraints between each modality and the predicted target. Chen et al. [6] adopt a similar approach, enhancing multimodal fusion by adding cross-attention. Furthermore, Liu et al. [28] apply IB during the representation fusion stage to discard irrelevant information from individual modalities. Xiao et al. [47] propose a two-layer IB structure that minimizes mutual information between input and latent states while maximizing it between latent and residual states. Moreover, conditional mutual information [46] has demonstrated strong generalization performance in various applications, such as feature selection [13], modality enhancement [23], and multimodal selection [20]. Gondek et al. [17] propose maximizing conditional mutual information to obtain relevant but novel clustering solutions, avoiding redundant information already captured by known structures or categories. Li et al. [25] demonstrate through CIB calculations that higher modal complementarity reduces robustness while noise impacts highly complementary modes more severely. Thus, recognizing the potential of CIB [33] for extracting relevant information in multimodal scenarios, we design a CIB-based method for multimodal fusion.
3 Task Analysis and Refactoring
@@
Heuristics in Character Role Labeling. Capturing character-specific traits can improve sarcasm detection, but methods relying on character association lack generalizability, as such features are often unavailable in real-world settings. We analyzed the proportion of sarcasm and non-sarcasm in the dialogues of different characters in the MUStARD++. The chi-square test is used to examine the relationship between character and sarcasm, with the null hypothesis stating no association. The test statistic of 166.7 and p-value of $3.89\times 10^{-27}$ provide strong evidence against the null hypothesis. The actors exhibited a distinct tendency towards either sarcastic or non-sarcastic behavior, suggesting that character traits strongly influence the judgment of sarcasm. The experimental findings [3, 26] reveal that the majority of models can enhance sarcasm recognition merely by incorporating character embeddings. However, the performance of such models trained with character embeddings will drop drastically when tested on other sentiment tasks. Thus, it is essential to design a robust model that captures the character’s intended features rather than relying on character labels.
@@
Label Leakage from Canned Laughter. Since most sarcasm datasets are derived from sitcoms, they typically include canned laughter. Canned laughter, a pre-recorded sound used to enhance humor or guide audience reactions, is also prevalent in MUStARD++. It often follows ironic or humorous utterances, leading to label leakage. Most studies [2, 42, 26] ignore the canned laughter while processing the data. Using the pre-trained speech recognition model SpeechPrompt v2[4], we validated on MUStARD++ and MUStARD++R (with and without canned laughter). Compared to MUStARD++, performance on MUStARD++R shows a significant drop, with F1-score decreasing from 73.47 to 43.59 (a drop of 29.88) and accuracy from 78.33 to 63.03 (a drop of 15.3). This suggests that the model learned the features associated with canned laughter, indirectly confirming the label leakage issue.
@@
Shortcuts of Emotional Inconsistency. Sarcasm is often a sugar-coated bomb, expressing the opposite of its literal meaning by masking implicit emotions with explicit ones. The statistical results from the MUStARD++ dataset reveal that sarcasm is predominantly associated with different emotions, accounting for $99\%$ of cases, while only $1\%$ is linked to the same emotions. In contrast, non-sarcasm is mainly associated with the same emotions ($94.7\%$), with only $5.3\%$ corresponding to different emotions. This indicates that sarcastic sentences often exhibit discrepancies between explicit and implicit sentiment labels. The Phi coefficient analysis supports this observation, revealing a very strong association between emotional consistency and sarcasm ($\phi=0.94$). Many studies [5, 36, 21] use this pattern to detect sarcasm. But in real-world scenarios, the explicit/implicit emotion labels do not exist. Moreover, inferring explicit/implicit emotion labels may be more challenging than directly detecting sarcasm. These limitations suggest that relying on explicit/implicit emotion labels for MSD is unfeasible.
@@ 4 Methodology
Modal Fusion Rethink. We rethink the limited effectiveness of previous fusion methods in light of the complementary and redundant properties of multimodal data. To provide quantitative evidence, we conduct ablation experiments on different modalities (see the appendix). Ablation tests on various combinations of modality reveal limited gains from modality fusion. In some cases, adding modalities reduces performance: audio may decrease overall effectiveness. Figure 3 illustrates the reasons for ineffective modality fusion: the Redundancy region is larger than the Complementarity region, and the overlapping middle region is too small. This suggests that misleading redundant information outweighs the informational gain from the added modality. For effective modality addition, irrelevant information that may mislead predictions should be reduced, while enhancing the benefits of complementary useful information.
@@ 4.1 Our Framework
where $L$ is the number of samples. Detailed derivations of the equations are elaborated in the appendix.
@@
To uncover latent information within each modality, we design a module for fine-grained feature extraction. Using GENTLE [39], we align audio and segment audio at the word level, which then serves as a reference for aligning fine-grained features in the visual modality. Next is the multimodal feature fusion module MCIB, which integrates complementary information across modalities while filtering out redundancy. Specifically, the MCIB minimizes the mutual information between the primary modality $x_{p}$ and the latent state $b$, condensing $x_{p}$ and filtering out redundant information. Simultaneously, it maximizes the conditional mutual information between the auxiliary modality $x_{a}$, the latent state $b$, and the target $y$, aiming to incorporate additional relevant information from $x_{a}$ so that $b$ holds useful information for predicting $y$. The model concatenates the trained latent state $b$ for prediction, which contains the ”redundancy-removed, effective complementarity” information distilled through the conditional information bottleneck.
@@
5 Experiments
Dataset and Evaluation Metrics. Experiments were conducted on MUStARD++ and MUStARD++R, respectively. Given the balanced class distribution, we report results using weighted precision, weighted recall, and weighted F1-score metrics. In addition, we tested the MCIB method on multimodal sentiment analysis tasks (CMU-MOSI and CMU-MOSEI in the appendix), achieving highly competitive results, which validate its generalization capability.
Implementation Details. To extract fine-grained features, we utilize several high-performance backbones. Text features are obtained using a pre-trained DeBERTa [19], yielding a representation size of $d_{t}=768$. For audio features, Mel Frequency Cepstral Coefficients (MFCC) and Mel spectrograms are generated with Librosa [32], along with prosodic features from OpenSMILE 3 [12], resulting in a combined representation of $d_{a}=291$. The video features are extracted by processing utterance frames through the pool5 layer of a ResNet-152 [18] pre-trained on ImageNet [8], producing a visual representation of size $d_{v}=2048$. All experiments were conducted on Nvidia A100 GPUs with 40GB of memory. Multiple trials demonstrate that combinations of random hyperparameters within $[1,64]$ typically result in local optima after 10 searches. Experimental results were averaged over five runs. Full statistics are provided in the appendix.
5.1 Comparison with baseline models
@@
We compare our proposed MCIB with the following baselines on multimodal fusion and sentiment analysis. VyAnG-Net [34] enhances sarcasm detection by integrating visual-specific attention mechanisms with text captions, while PredGaze [40] improves sarcasm recognition by utilizing fine-grained visual information such as eye gaze. FIES [14] proposes a multimodal approach that integrates audio, textual, sentiment, and emotion data to enhance sarcasm detection, while ABCA-IMI [26] identifies sarcasm through multiple inconsistency detection mechanisms. Some methods did not release their code; for instance, the VyAnG-Net results were obtained by training on MUStARD and validating on MUStARD++. SIB [31, 6] using the IB between each single modality and the target, while DIB [52] applies the IB to pairs of modalities, enabling back-optimization. ITHP [47] was designed with a two-layer IB guide to the modality information flow. TBJE [7] is a cross-attention-based model with high generalization for multimodal fusion. MUStARD method [35] leverages a collaborative gating mechanism for sarcasm and emotion recognition. Moreover, we report baseline results for the pre-trained speech model SpeechPrompt v2 [4], and leading large language models: OpenAI’s latest flagship model GPT-4o [11], the newest release from Google DeepMind Gemini 2.5 [10], under multimodal configurations.
@@
Additionally, we conducted cross-training experiments between MUStARD++ and MUStARD++R to further investigate the impact of shortcut learning on multimodal sarcasm detection (see Appendix).
5.2 Modality and Module Ablation Studies
@@ 5.3 Visualizations
Figure 6 provides an intuitive comparison of the effectiveness of the proposed MCIB method in modality fusion relative to baseline models. The directional trends of the arrows indicate that MCIB optimizes in the correct direction for modality fusion. Observing the overall improvement: misclassifications due to redundant or irrelevant information are significantly reduced across all modality combinations, while complementary information between modalities is effectively utilized. The accuracy of joint predictions from added modalities has also improved. Further examples are provided in the appendix.
6 Conclusion

### v2

@@ Abstract
Codehttps://github.com/sljgkjhwe/MCIB.git
Datasetshttps://pan.quark.cn/s/975e0d976744#/list/share (3BFf)
Introduction
Multimodal sentiment analysis (MSA) (zadeh2017tensor) integrates various modalities to interpret human emotions, sentiments, and opinions. In MSA tasks, multimodal sarcasm detection (castro2019towards) is particularly challenging due to the subtle contrast between surface meaning and underlying intent, which may convey humor, ridicule, or contempt (joshi2017automatic). See Figure 1, Sheldon’s words express surprise, but his facial expression shows disgust, and his tone is mocking, conveying sarcasm. Despite advances in large language models, their performance remains capped by supervised baselines, falling short of human-level understanding (hu2025emobenchmbench).
@@ Introduction
Numerous studies have explored multimodal sarcasm detection (MSD) from various perspectives. Many focus on the inherent characteristics of sarcasm or are specifically designed for prior information: such as contrasting emotional polarity (lu2024fact; wang2024cross), or incorporate external knowledge (yue2023knowlenet), character profiles (li2024attention), emotional cues (gao2024improving), and fine-grained visual features (tiwari2023predict).
As shown by issues 1, 2, and 3 in Figure 1, models tend to learn specific shortcuts. Those shortcuts (geirhos2020shortcut) often fail to generalize to more diverse conditions, typically in real-world scenarios. We identify three major shortcut issues in multimodal sarcasm detection research with the representative dataset MUStARD++ (ray2022multimodal). (1) Heuristics in Character Role Labeling: In sarcasm datasets, some characters exhibit a speaking style that favors sarcasm, such as Sheldon in the TV show THE BIG BANG THEORY. Introducing these character labels during training causes the model to form a bias toward sarcasm associated with a specific person. (2) Label Leakage from Canned Laughter: Canned laughter (artificial background laughter) often causes label leakage. A review of MUStARD++ reveals that canned laughter frequently follows sarcastic utterances, but rarely non-sarcastic ones. Thus, the model tends to classify samples with canned laughter as sarcasm. (3) Shortcuts of Emotional Inconsistency: The dataset provides implicit/explicit emotion labels, and their contrast serves as a strong indicator of sarcasm. Some models are designed to learn implicit/explicit emotions rather than true sarcasm, limiting their ability to detect sarcasm due to over-reliance on inconsistency.
The key to understanding the intended complex information and genuinely improving performance is addressing the core multimodal fusion problem, rather than designing modules based on shortcut learning. Various fusion methods (ray2022multimodal; zhang2024self; zhang2023learning) do not result in significant information gains, some even reducing task performance, highlighting their poor generalization and adaptability to sarcasm tasks. Higher modal complementarity leads to poorer robustness when certain modes are absent or misaligned, and noise has a greater impact on modes with high complementarity (li2023modality). Multimodal sarcasm detection involves heterogeneous information with complementary but also irrelevant or misleading redundancy. Therefore, this underscores the need for an efficient fusion strategy that integrates complementary information while mitigating redundancy.
@@
Multimodal Sarcasm Detection has evolved from text and image-text (wen2023dip) to video clips, which better reflect real-life scenarios. Many methods leverage sarcasm traits to aid recognition, such as incongruity (lu2024fact). Song et al. (song2024utterance) propose an utterance-level attention and incongruity learning network to capture incongruity representations in sarcastic expressions. Since sarcasm is closely linked to emotions and sentiments, an adaptive representation learning model (zhang2024self) is proposed based on an expert network to analyze the emotion and sarcasm tasks jointly. Chauhan et al. (chauhan2020sentiment) design a multitasking collaborative framework for cross-training and sharing attention between sarcasm, implicit/explicit emotions, and implicit/explicit sentiments. Additionally, Zhang et al. (zhang2023learning) employ a dual-gating network and three separate layers to achieve cross-modal interactive learning. Other research attempts to leverage various external information. Studies such as (liang2022multi) and (yue2023knowlenet) have utilized tools like KnowleNet to implement knowledge fusion networks for MSD. Moreover, works like (tomar2024action), (tiwari2023predict), and (pandey2025vyang) enhance sarcasm recognition by incorporating conversational behavioral cues, fine-grained eye gaze information, or deep visual features acquired through self-conditioning ConvNet, respectively. While (gao2024improving; helal2024contextual) rely on pre-training with explicit and implicit emotions, the best results from (li2024attention) similarly through character labeling. Besides, some novel approaches (liu2023quantum; tiwari2024quantum) suggest that quantum mechanics are well-suited for capturing the complexity and uncertainty in sarcasm through a quantum probability-driven model or quantum fuzzy neural networks.
@@
Information Bottleneck (IB) based approaches (slonim1999agglomerative; alemi2022deep) have demonstrated their effectiveness in various tasks such as cross-modal clustering (yan2023cross) and representation learning (zhang2022unimodal; ding2023robust), and have also been widely applied in multimodal fusion. For example, Zhang et al. (zhang2022information) improve multimodal sentiment analysis performance by applying IB constraints to pairs of modalities. Mai et al. (mai2022multimodal) ensure that latent modal representations can effectively handle the target task by introducing IB constraints between each modality and the predicted target. Chen et al. (chen2023multimodal) adopt a similar approach, enhancing multimodal fusion by adding cross-attention. Furthermore, Liu et al. (liu2024multimodal) apply IB during the representation fusion stage to discard irrelevant information from individual modalities. Xiao et al. (xiao2024neuro) propose a two-layer IB structure that minimizes mutual information between input and latent states while maximizing it between latent and residual states. Moreover, conditional mutual information (wyner1978definition) has demonstrated strong generalization performance in various applications, such as feature selection (fleuret2004fast), modality enhancement (ji2022increasing), and multimodal selection (he2024efficient). Gondek et al. (gondek2003conditional) propose maximizing conditional mutual information to obtain relevant but novel clustering solutions, avoiding redundant information already captured by known structures or categories. Li et al. (li2023modality) demonstrate through CIB calculations that higher modal complementarity reduces robustness while noise impacts highly complementary modes more severely. Thus, recognizing the potential of CIB (molavipour2021neural) for extracting relevant information in multimodal scenarios, we design a CIB-based method for multimodal fusion.
Task Analysis and Refactoring
@@
Heuristics in Character Role Labeling. Capturing character-specific traits can improve sarcasm detection, but methods relying on character association lack generalizability, as such features are often unavailable in real-world settings. We analyzed the proportion of sarcasm and non-sarcasm in the dialogues of different characters in the MUStARD++. The chi-square test is used to examine the relationship between character and sarcasm, with the null hypothesis stating no association. The test statistic of 166.7 and p-value of $3.89\times 10^{-27}$ provide strong evidence against the null hypothesis. The actors exhibited a distinct tendency towards either sarcastic or non-sarcastic behavior, suggesting that character traits strongly influence the judgment of sarcasm. The experimental findings (castro2019towards; li2024attention) reveal that the majority of models can enhance sarcasm recognition merely by incorporating character embeddings. However, the performance of such models trained with character embeddings will drop drastically when tested on other sentiment tasks. Thus, it is essential to design a robust model that captures the character’s intended features rather than relying on character labels.
@@
Label Leakage from Canned Laughter. Since most sarcasm datasets are derived from sitcoms, they typically include canned laughter. Canned laughter, a pre-recorded sound used to enhance humor or guide audience reactions, is also prevalent in MUStARD++. It often follows ironic or humorous utterances, leading to label leakage. Most studies (arora2023universlu; tomar2023your; li2024attention) ignore the canned laughter while processing the data. Using the pre-trained speech recognition model SpeechPrompt v2(chang2023speechprompt), we validated on MUStARD++ and MUStARD++R (with and without canned laughter). Compared to MUStARD++, performance on MUStARD++R shows a significant drop, with F1-score decreasing from 73.47 to 43.59 (a drop of 29.88) and accuracy from 78.33 to 63.03 (a drop of 15.3). This suggests that the model learned the features associated with canned laughter, indirectly confirming the label leakage issue.
@@
Shortcuts of Emotional Inconsistency. Sarcasm is often a sugar-coated bomb, expressing the opposite of its literal meaning by masking implicit emotions with explicit ones. The statistical results from the MUStARD++ dataset reveal that sarcasm is predominantly associated with different emotions, accounting for $99\%$ of cases, while only $1\%$ is linked to the same emotions. In contrast, non-sarcasm is mainly associated with the same emotions ($94.7\%$), with only $5.3\%$ corresponding to different emotions. This indicates that sarcastic sentences often exhibit discrepancies between explicit and implicit sentiment labels. The Phi coefficient analysis supports this observation, revealing a very strong association between emotional consistency and sarcasm ($\phi=0.94$). Many studies (chauhan2020sentiment; shah2022emotion; helal2024contextual) use this pattern to detect sarcasm. But in real-world scenarios, the explicit/implicit emotion labels do not exist. Moreover, inferring explicit/implicit emotion labels may be more challenging than directly detecting sarcasm. These limitations suggest that relying on explicit/implicit emotion labels for MSD is unfeasible.
@@
Modal Fusion Rethink. We rethink the limited effectiveness of previous fusion methods in light of the complementary and redundant properties of multimodal data. Ablation tests on various combinations of modality reveal limited gains from modality fusion. In some cases, adding modalities reduces performance: audio may decrease overall effectiveness. Figure 3 illustrates the reasons for ineffective modality fusion: the Redundancy region (which refers not only to ”repeated information” that may be useless for the object but also to negative interference in modality fusion) is larger than the Complementarity region, and the overlapping middle region is too small. This suggests that misleading redundant information outweighs the informational gain from the added modality. For effective modality addition, irrelevant information that may mislead predictions should be reduced, while enhancing the benefits of complementary useful information.
@@
where $L$ is the number of samples.
@@
To uncover latent information within each modality, we design a module for fine-grained feature extraction. Using GENTLE, we align audio and segment audio at the word level, which then serves as a reference for aligning fine-grained features in the visual modality. Next is the multimodal feature fusion module MCIB, which integrates complementary information across modalities while filtering out redundancy. Specifically, the MCIB minimizes the mutual information between the primary modality $x_{p}$ and the latent state $b$, condensing $x_{p}$ and filtering out redundant information. Simultaneously, it maximizes the conditional mutual information between the auxiliary modality $x_{a}$, the latent state $b$, and the target $y$, aiming to incorporate additional relevant information from $x_{a}$ so that $b$ holds useful information for predicting $y$. The model concatenates the trained latent state $b$ for prediction, which contains the ”redundancy-removed, effective complementarity” information distilled through the conditional information bottleneck.
@@
Experiments
Dataset and Evaluation Metrics. Experiments were conducted on MUStARD++ and MUStARD++R, respectively. Given the balanced class distribution, we report results using weighted precision, weighted recall, and weighted F1-score metrics. In addition, we tested the MCIB method on multimodal sentiment analysis tasks (CMU-MOSI and CMU-MOSEI in the Extended version), achieving highly competitive results, which validate its generalization capability.
Implementation Details. To extract fine-grained features, we utilize several high-performance backbones. Text features are obtained using a pre-trained DeBERTa (he2021debertav3), yielding a representation size of $d_{t}=768$. For audio features, Mel Frequency Cepstral Coefficients (MFCC) and Mel spectrograms are generated with Librosa (McFee2018librosalibrosa0), along with prosodic features from OpenSMILE 3 (eyben2010opensmile), resulting in a combined representation of $d_{a}=291$. The video features are extracted by processing utterance frames through the pool5 layer of a ResNet-152 (he2016deep) pre-trained on ImageNet (deng2009imagenet), producing a visual representation of size $d_{v}=2048$. All experiments were conducted on Nvidia A100 GPUs with 40GB of memory. Multiple trials demonstrate that combinations of random hyperparameters within $[1,64]$ typically result in local optima after 10 searches. The hyperparameter analysis reveals that the most significant interactions occur between $\lambda_{2}\And\alpha_{1}$, $\lambda_{1}\And\alpha_{1}$, and $\alpha_{0}\And\alpha_{1}$. The stability analysis highlighted that $\alpha_{1}$ demonstrated the best stability. Experimental results were averaged over five runs.
Comparison with baseline models
@@
We compare our proposed MCIB with the following baselines on multimodal fusion and sentiment analysis. VyAnG-Net (pandey2025vyang) enhances sarcasm detection by integrating visual-specific attention mechanisms with text captions, while PredGaze (tiwari2023predict) improves sarcasm recognition by utilizing fine-grained visual information such as eye gaze. FIES (gao2024improving) proposes a multimodal approach that integrates audio, textual, sentiment, and emotion data to enhance sarcasm detection, while ABCA-IMI (li2024attention) identifies sarcasm through multiple inconsistency detection mechanisms. Some methods did not release their code; for instance, the VyAnG-Net results were obtained by training on MUStARD and validating on MUStARD++. SIB (mai2022multimodal; chen2023multimodal) using the IB between each single modality and the target, while DIB (zhang2022information) applies the IB to pairs of modalities, enabling back-optimization. ITHP (xiao2024neuro) was designed with a two-layer IB guide to the modality information flow. TBJE (delbrouck2020transformer) is a cross-attention-based model with high generalization for multimodal fusion. MUStARD method (ray2022multimodal) leverages a collaborative gating mechanism for sarcasm and emotion recognition. Moreover, we report baseline results for the pre-trained speech model SpeechPrompt v2 (chang2023speechprompt), and leading large language models: OpenAI’s latest flagship model GPT-4o (openai2024gpt4ocard), the newest release from Google DeepMind Gemini 2.5 (comanici2025gemini25pushingfrontier), under multimodal configurations.
@@
Additionally, we conducted cross-training experiments between MUStARD++ and MUStARD++R to further investigate the impact of shortcut learning on multimodal sarcasm detection (see Extended version).
Modality and Module Ablation Studies
@@
Figure 6 provides an intuitive comparison of the effectiveness of the proposed MCIB method in modality fusion relative to baseline models. The directional trends of the arrows indicate that MCIB optimizes in the correct direction for modality fusion. Observing the overall improvement: misclassifications due to redundant or irrelevant information are significantly reduced across all modality combinations, while complementary information between modalities is effectively utilized. The accuracy of joint predictions from added modalities has also improved.
Conclusion